OpenStudy (anonymous):

VERY DIFFICULT WORD PROBLEM. A Bike has 1 Handlebar, 2 Pedals and 2 Wheels. A Tricycle has 1 Handlebar, 2 Pedals, and 3 Wheels. A Tandem Bike has 1 Handlebar, 4 Pedals, and 2 Wheels. If a Bikeshop has a total of 144 Handlebars, 378 Pedals and 320 Wheels, how many Bicycles, Tandem Bikes, and Tricycles does it have? Please at show work, or explain how to solve this.

OpenStudy (blockcolder):

Let B be the number of bikes, T the number of trikes, and D the number of tandems. There are B+T+D=144 handlebars, 2B+2T+4D=378 pedals, and 2B+3T+2D=320 wheels. Solve the system to get your answer. :)

OpenStudy (anonymous):

I'm not too sure what to do from that point on...

OpenStudy (blockcolder):

Here, I'll start: \[\begin{cases} B+T+D=144 \qquad &1\\ 2B+2T+4D=378 &2\\ 2B+3T+2D=320 &3 \end{cases}\] Divide equation 2 by 2: \[B+T+2D=189\] Subtract equation 1 from this: \[D=189-144=45\] Therefore, from the very first equation: \[B+T=144-45=99\] Now subtract equation 2 from 3: \[T-2D=-58\\ T=-58+2D=32\] Can you determine what B is now? :)

OpenStudy (anonymous):

...I think I get it... ...Is B... 67?

OpenStudy (blockcolder):

Right-o! :)

OpenStudy (anonymous):

Then T must be 32, and D must be 45! Thank you!

OpenStudy (blockcolder):

You're welcome. :D